OpenStudy (anonymous):

-14x=-7-7Y -10X=-15-7Y

OpenStudy (texaschic101):

lets rearrange.. -14x = -7 - 7y -14x + 7 = -7y -14x/-7 + 7/-7 = -7y/-7 2x - 1 = y so now we sub 2x - 1 in for y in the other equation to find x -10x = -15 - 7y -10x = -15 - 7(2x - 1) --- distribute through the parenthesis -10x = -15 - 14x + 7 --- add 14x to both sides -10x + 14x = -15 + 7 --- combine like terms 4x = - 8 x = -8/4 x = -2 now sub -2 in for x in either of the original equations to find y -14x = -7 - 7y -14(-2) = -7 - 7y 28 = -7 - 7y 28 + 7 = -7y 35 = -7y 35/-7 = y -5 = y check.. -10x = -15 - 7y -10(-2) = -15 - 7(-5) 20 = -15 + 35 20 = 20 (correct) so x = -2 and y = -5